# CastMind: An Interaction-Driven Agentic Reasoning Framework for Cognition-Inspired Time Series Forecasting

### v2

@@ Abstract
Time series forecasting plays a crucial role in decision-making across many real-world applications. Despite substantial progress, most existing methods still treat forecasting as a static, single-pass regression problem. In contrast, human experts form predictions through iterative reasoning that integrates temporal features, domain knowledge, case-based references, and supplementary context, with continuous refinement. In this work, we propose CastMind, an interaction-driven agentic reasoning framework that enables accurate time series forecasting with training-free large language models. CastMind reformulates forecasting as an expert-like process and organizes it into a multi-stage workflow involving context preparation, reasoning-based generation, and reflective evaluation, transforming forecasting from a single-pass output into a multi-turn, autonomous interaction process. To support diverse perspectives commonly considered by human experts, we develop a lightweight toolkit comprising a feature set, a knowledge base, a case library, and a contextual pool that provides external support for LLM-based reasoning. Extensive experiments across multiple benchmarks show that CastMind generally outperforms representative baselines [^1] https://github.com/SkyeGT/CastMind.
@@ 1 Introduction
Figure 1: Illustration of the motivation behind CastMind: shifting forecasting from static, single-pass regression to human-like, interactive, multi-turn reasoning.
Time series forecasting plays a crucial role in decision-making across various real-world applications, such as energy (Cao et al., 2023), climate (Wang et al., ), and industry (Ma et al., 2025). Despite advances with statistical (Webby & O’Connor, 1996), deep learning (Benidis et al., 2022), and recent foundation models (Das et al., 2024), these methods typically treat forecasting as a static, single-pass regression task, generating predictions based solely on historical observations (Brockwell & Davis, 2002). This approach neglects reasoning, interaction, and iterative refinement, which are critical in real-world decision-making.
@@
To address this, we argue that time series forecasting be reformulated as an interaction-driven agentic reasoning problem. Rather than a single-pass prediction, forecasting becomes a multi-turn process where an agent actively prepares contextual information, reasons over diverse sources, and refines predictions through self-reflection. To operationalize this perspective, We introduce CastMind, an interaction-driven agentic reasoning framework for time series forecasting. CastMind leverages training-free LLMs as reasoning engines and organizes forecasting into a structured workflow: context preparation, reasoning-based generation, and reflective verification. By modeling interaction and refinement, CastMind bridges the gap between existing models and expert-like forecasting processes.
CastMind implements this through a congnitive-inspired workflow that separates reasoning from prediction. It starts with context preparation, gathering relevant information beyond raw time series, including temporal features, domain knowledge, case-based references and supplementary context. The LLM then analyzes this context and produces a intermediate forecast through reasoning. A reflective verification stage follows, ensuring consistency and plausibility, enabling self-correction when necessary. To support this process, CastMind uses a lightweight toolkit with core components: a feature set, knowledge base, case library,and contextual pool. Together, the workflow and toolkit enable CastMind to conduct forecasting as an interactive, multi-turn reasoning process that closely mirrors expert practice.
@@
  • We propose CastMind, a training-free agentic reasoning framework that integrates a structured workflow with a lightweight forecasting toolkit for context-aware and reflective forecasting.
  • Extensive experiments across multiple benchmarks show CastMind consistently outperforms statistical, deep learning, and foundation-model baselines, demonstrating its effectiveness.
@@ 2.2 LLM-based Reasoning
Recent advancements in LLMs have demonstrated remarkable generalization capabilities in reasoning and decision-making (Guo et al., 2025). Foundational is in-context learning (ICL), enabling models to adapt to unseen tasks via prompts without parameter updates (Brown et al., 2020; Xue & Salim, 2023). To elicit multi-step inference, prompting strategies like chain-of-thought (CoT) (Wei et al., 2022) decompose complex problems into intermediate reasoning steps. Subsequent research enhanced robustness through aggregation like self-consistency (Wang et al., 2022) and decomposition like least-to-most prompting (Zhou et al., 2022). Moving beyond static reasoning, a significant paradigm shift models LLMs as autonomous agents operating in multi-step, goal-directed settings. Unlike passive inference, agentic systems maintain intermediate states and interleave reasoning with actions, invoking tools to ground outputs. Frameworks like ReAct (Yao et al., 2023), Toolformer (Schick et al., 2023), and WebGPT (Nakano et al., 2021) exemplify this approach by integrating search engines and APIs to improve reliability. To handle increasingly complex tasks, recent agents evolved into comprehensive cognitive architectures. These systems extend tool use by incorporating modules like structured planning (Wang et al., 2023; Erdogan et al., 2025), long-term memory (Packer et al., 2023), and self-correction like reflexion (Shinn et al., 2023). This modular design enables refinement and adaptive execution, forming the basis for our framework.
Figure 2: Overall architecture of CastMind. The three-stage process (Investigator, Generator, Reflector) bridges human wisdom and LLM intelligence, simulating expert cognition to transform forecasting into an interactive, multi-turn reasoning process.
3 The Proposed CastMind
In this section, we introduce CastMind, which reformulates forecasting as a multi-stage process that simulates human-like reasoning to prepare, generate, and evaluate predictions.
@@ 3.2 Overall Agentic Workflow
As shown in Figure 2, CastMind is a multi-stage interaction-driven agentic reasoning framework designed for cognition-inspired time series forecasting. It reformulates forecasting from a static, single-pass mapping task into an interactive, multi-turn iterative process with training-free LLM reasoning. Before the workflow begins, CastMind prepares a specialized toolkit comprising a feature set, a knowledge base, a case library, a contextual pool. The process then unfolds in three coordinated stages. Initially, CastMind organizes useful contextual input and information retrieved from the toolkit to construct a well-defined forecasting context. Subsequently, the framework operates on this foundation to generate reasoning-based predictions. Finally, it evaluates intermediate outputs and the reason trace to detect inconsistencies and iteratively refine the predictions. Overall, this workflow mirrors how human experts extract context, reason about predictions, and adjust decisions.
@@ 3.3 Forecasting Toolkit Design
CastMind extracts temporal features from endogenous time series and exogenous variables to summarize temporal dynamics. Given a group of extraction functions $\mathcal{F}=\{f_{1},\dots,f_{n}\}$, the feature set $F$ is constructed by applying each function to all input series:
@@
CastMind maintains a knowledge base that provides prior information to support reasoning and decision-making. The knowledge base contains: (1) practical rules, which are experience-based guidelines derived from empirical observations. These rules capture common patterns in the data, such as the effect of rainfall on water levels or the correlation between hot weather and increased electricity demand. (2) theoretical principles, which are domain-specific laws and theorems that help constrain the model’s predictions. These include well-established principles such as Joule’s law and the conservation of mass, which ensure that the model’s predictions align with physical and scientific laws. Let KKdenote the domain knowledge retrieved from the knowledge base.
@@
To simulate the expert use of historical experiences, CastMind constructs a case library by associating each training sample $\langle X_{i},Y_{i}\rangle$ with its optimal candidate model $M_{i}$ selected from the candidate pool $\mathcal{M}$ (detailed in Appendix B.2). To enable efficient retrieval and generation of case-based predictions, we perform K-means clustering (detailed in Appendix A.2) on all $X_{i}$, grouping them into multiple clusters. Each cluster $C_{m}$ contains training samples $X_{i}$ that are most similar to one another. The center $c_{m}$ of each cluster $C_{m}$ is determined by averaging all the samples within the cluster, where the center is the mean of the samples $X_{i}$ in $C_{m}$, calculated as $c_{m}=\frac{1}{|C_{m}|}\sum_{X_{i}\in C_{m}}X_{i}$.
@@ 3.4 The Investigator: Context Preparation
Accurate forecasting requires a well-defined prediction task, a proper understanding of data semantics, and the availability of relevant contextual information. Therefore, CastMind constructs a standardized contextual input $I_{\text{in}}$ that integrates time series with textual descriptions. Specifically, $I_{\text{in}}$ comprises the task prompt, the data profile (detailed meanings of the variables), original timestamp information (Wang et al., 2024a; Zeng et al., 2024), and the raw time series $(X_{\text{ex}},X_{\text{en}})$ in original space. This standardized input serves as the foundation for subsequent analysis.
@@ 3.5 The Generator: Reasoning-based Forecasting
Given the consolidated context $I_{\text{con}}$, CastMind derives an intermediate prediction $Y_{\text{int}}$ through reasoning-driven generation instead of single-pass regression. The model initializes with the case-based prediction $Y_{\text{case}}$ as a reliable reference. To handle distribution shifts between neighbor $X_{\text{nb}}$ and the current sequence $X_{\text{en}}$, CastMind uses the neighbor trajectory $Y_{\text{nb}}$ to adjust for scale and fluctuation differences (Liu et al., 2025c). The framework also incorporates temporal features from the selected feature set $F_{\text{sel}}$, along with domain knowledge $K$ and supplementary context $S$. These elements support more sophisticated adjustments, such as shifting or reshaping, that allow the model to adapt to subtle but important changes in the data, which may be overlooked in simpler models. By incorporating timestamps $T$ for time-aware calibration (e.g., distinguishing weekday variations), the reasoning process is formalized as $Y_{\text{int}}=\mathcal{G}(I_{\text{con}})$, where $\mathcal{G}$ denotes the generator that integrates this context to output the adjusted $Y_{\text{int}}$ for the subsequent stage.
@@ 3.6 The Reflector: Iterative Refinement
In real-world forecasting, predictions are usually reviewed and refined before being finalized. Accordingly, CastMind includes a reflection stage to verify the prediction and support iterative refinement. This stage improves the overall reliability and reproducibility of the framework and consists of two parts: (1) prediction evaluation, which assesses whether the prediction satisfies the task requirements and is supported by the available evidence, ensuring that the prediction is reasonable and reliable; and (2) reasoning-trace evaluation, which examines the generated reasoning trace to ensure that each step is valid and consistent, and identifies potential issues such as unsupported claims or logical gaps. Together, these checks help ensure that both the forecast and its underlying reasoning adhere to the task requirements. If issues are found, CastMind applies targeted, evidence-supported corrections to $Y_{\text{int}}$ to obtain a refined prediction. The refined result is then re-evaluated with the same criteria. If acceptable, CastMind outputs the forecast; otherwise, it returns feedback to the previous stage for another update, repeating until an acceptable result is obtained.
@@ 4 Experiments
In this section, we conduct experiments on various forecasting benchmarks comparing CastMind with mainstream approaches to demonstrate its effectiveness.
@@ 4.2 Main Results
Table 1 compares CastMind against twelve baselines across short-term and long-term benchmarks. Overall, CastMind achieves state-of-the-art performance with superior accuracy across most metrics. In short-term scenarios, CastMind shows outstanding stability in the face of high volatility, effectively managing unpredictable fluctuations. Unlike statistical methods such as ARIMA, which struggle with rapid non-linear fluctuations and inherent stochastic volatility in price dynamics, or deep learning models such as PatchTST, which may over-smooth dynamics, CastMind effectively captures sharp peaks by leveraging historical references to avoid phase shifts. This gives it a significant advantage on datasets like BE and FR. In long-term settings, CastMind exhibits superior forecasting capability, especially on datasets like Windy Power and Sunny Power, which exhibit complex environmental dependencies. While foundation models such as Chronos perform competitively due to their zero-shot generalization, CastMind maintains a consistent lead through its reasoning and refinement capabilities that enforce causal constraints derived from diverse contexts. For instance, on the Windy Power dataset, CastMind substantially outperforms baselines like DLinear. This superiority across different scales validates the effectiveness of the interactive, multi-turn reasoning process in handling diverse temporal patterns and ensuring forecasting stability.
@@ 4.3 Ablation studies
This section validates the CastMind framework through a multi-perspective ablation study, covering both architectural stages, contextual input, and auxiliary toolkit modules.
@@
To validate the workflow, we conduct an ablation study by removing the toolkit, reasoning, and reflexion. As shown in Table 2, the full model consistently achieves best performance, validating component collaboration. Pecifically, removing the toolkit yields highest errors in scenarios like BE and ETTm, confirming external context is foundational. Furthermore, comparing the baseline lacking reasoning, which serves as the state-of-the-art candidate model backbone, with the full model reveals that reasoning-based forecasting substantially improves accuracy, evidenced by notable error reduction on the FR dataset. Most importantly, results on the Sunny Power dataset highlight the indispensability of the reflexion module, where the variant excluding self-correction performs worse than even the non-reasoning baseline. This anomaly suggests that without self-correction, generative reasoning may be prone to hallucinations or overfitting, an issue rectified by the full model acting as a safeguard for reliability and physical consistency.
@@
To verify the necessity of constructing a comprehensive context $I_{\text{con}}$, we conduct a component-wise ablation study. As shown in Figure 3, the complete CastMind framework consistently yields the lowest MAE, confirming the synergistic contribution of all elements. Specifically, removing timestamps causes significant degradation on periodic datasets such as PJM and Sunny Power, highlighting the critical role of explicit temporal identifiers in capturing seasonality that numerical sequences alone cannot convey. Similarly, the exclusion of exogenous variables and static data profile leads to substantial error increases on tasks like ETTm and NP, demonstrating that external drivers and domain-specific inductive biases are essential for reasoning about complex system behaviors beyond historical autoregressive patterns.

### v3

# AlphaCast: An Interaction-Driven Agentic Reasoning Framework for Cognition-Inspired Time Series Forecasting

@@ Abstract
Time series forecasting plays a crucial role in decision-making across many real-world applications. Despite substantial progress, most existing methods still treat forecasting as a static, single-pass regression problem. In contrast, human experts form predictions through iterative reasoning that integrates temporal features, domain knowledge, case-based references, and supplementary context, with continuous refinement. In this work, we propose AlphaCast, an interaction-driven agentic reasoning framework that enables accurate time series forecasting with training-free large language models. AlphaCast reformulates forecasting as an expert-like process and organizes it into a multi-stage workflow involving context preparation, reasoning-based generation, and reflective evaluation, transforming forecasting from a single-pass output into a multi-turn, autonomous interaction process. To support diverse perspectives commonly considered by human experts, we develop a lightweight toolkit comprising a feature set, a knowledge base, a case library, and a contextual pool that provides external support for LLM-based reasoning. Extensive experiments across multiple benchmarks show that AlphaCast generally outperforms representative baselines [^1] https://github.com/SkyeGT/CastMind.
@@ 1 Introduction
Figure 1: Illustration of the motivation behind AlphaCast: shifting forecasting from static, single-pass regression to human-like, interactive, multi-turn reasoning.
Time series forecasting plays a crucial role in decision-making across various real-world applications, such as energy (Cao et al., 2023), climate (Wang et al., ), and industry (Ma et al., 2025). Despite substantial advances with statistical (Webby & O’Connor, 1996), deep learning (Benidis et al., 2022), and recent foundation models (Das et al., 2024), these methods typically treat forecasting as a static, single-pass regression task, generating predictions based solely on historical observations (Brockwell & Davis, 2002). This approach neglects reasoning, interaction, and iterative refinement, which are critical in complex real-world decision-making scenarios (Cheng et al., 2026).
@@
To address this, we argue that time series forecasting be reformulated as an interaction-driven agentic reasoning problem. Rather than a single-pass prediction, forecasting becomes a multi-turn process where an agent actively prepares contextual information, reasons over diverse sources, and refines predictions through self-reflection. To operationalize this perspective, We introduce AlphaCast, an interaction-driven agentic reasoning framework for time series forecasting. AlphaCast leverages training-free LLMs as reasoning engines and organizes forecasting into a structured workflow: context preparation, reasoning-based generation, and reflective verification. By modeling interaction and refinement, AlphaCast bridges the gap between existing models and expert-like forecasting processes.
AlphaCast implements this through a congnitive-inspired workflow that separates reasoning from prediction. It starts with context preparation, gathering relevant information beyond raw time series, including temporal features, domain knowledge, case-based references and supplementary context. The LLM then analyzes this context and produces a intermediate forecast through reasoning. A reflective verification stage follows, ensuring consistency and plausibility, enabling self-correction when necessary. To support this process, AlphaCast uses a lightweight toolkit with core components: a feature set, knowledge base, case library,and contextual pool. Together, the workflow and toolkit enable AlphaCast to conduct forecasting as an interactive, multi-turn reasoning process that closely mirrors expert practice.
@@
  • We propose AlphaCast, a training-free agentic reasoning framework that integrates a structured workflow with a lightweight forecasting toolkit for context-aware and reflective forecasting.
  • Extensive experiments across multiple benchmarks show AlphaCast consistently outperforms statistical, deep learning, and foundation-model baselines, demonstrating its effectiveness.
@@ 2.2 LLM-based Reasoning
Recent advancements in LLMs have demonstrated remarkable generalization capabilities in reasoning and decision-making (Guo et al., 2025). Foundational is in-context learning (ICL), enabling models to adapt to unseen tasks via prompts without parameter updates (Xue & Salim, 2023). To elicit multi-step inference, prompting strategies like chain-of-thought (CoT) (Wei et al., 2022) decompose complex problems into intermediate reasoning steps. Subsequent research enhanced robustness through aggregation like self-consistency (Wang et al., 2022) and decomposition like least-to-most prompting (Zhou et al., 2022). Moving beyond static reasoning, a significant paradigm shift models LLMs as autonomous agents operating in multi-step, goal-directed settings (Tao et al., 2026). Unlike passive inference, agentic systems maintain intermediate states and interleave reasoning with actions, invoking tools to ground outputs. Frameworks like ReAct (Yao et al., 2023), Toolformer (Schick et al., 2023), and WebGPT (Nakano et al., 2021) exemplify this approach by integrating search engines and APIs to improve reliability. To handle increasingly complex tasks, recent agents evolved into comprehensive cognitive architectures. These systems extend tool use by incorporating modules like structured planning (Wang et al., 2023; Erdogan et al., 2025), long-term memory (Packer et al., 2023), and self-correction like reflexion (Shinn et al., 2023). This modular design enables refinement and adaptive execution, forming the basis for our framework.
Figure 2: Overall architecture of AlphaCast. The three-stage process (Investigator, Generator, Reflector) bridges human wisdom and LLM intelligence, simulating expert cognition to transform forecasting into an interactive, multi-turn reasoning process.
3 The Proposed AlphaCast
In this section, we introduce AlphaCast, which reformulates forecasting as a multi-stage process that simulates human-like reasoning to prepare, generate, and evaluate predictions.
@@ 3.2 Overall Agentic Workflow
As shown in Figure 2, AlphaCast is a multi-stage interaction-driven agentic reasoning framework designed for cognition-inspired time series forecasting. It reformulates forecasting from a static, single-pass mapping task into an interactive, multi-turn iterative process with training-free LLM reasoning. Before the workflow begins, AlphaCast prepares a specialized toolkit comprising a feature set, a knowledge base, a case library, and a contextual pool. The process then unfolds in three coordinated stages. Initially, AlphaCast organizes useful contextual input and information retrieved from the toolkit to construct a well-defined forecasting context. Subsequently, the framework operates on this foundation to generate reasoning-based predictions. Finally, it evaluates intermediate outputs and the reason trace to detect inconsistencies and iteratively refine the predictions. Overall, this workflow mirrors how human experts extract context, reason about predictions, and adjust decisions.
@@ 3.3 Forecasting Toolkit Design
AlphaCast extracts temporal features from endogenous time series and exogenous variables to summarize temporal dynamics. Given a group of extraction functions $\mathcal{F}=\{f_{1},\dots,f_{n}\}$, the feature set $F$ is constructed by applying each function to all input series:
@@
AlphaCast maintains a knowledge base that provides prior information to support reasoning and decision-making. The knowledge base contains: (1) practical rules, which are experience-based guidelines derived from empirical observations. These rules capture common patterns in the data, such as the effect of rainfall on water levels or the correlation between hot weather and increased electricity demand. (2) theoretical principles, which are domain-specific laws and theorems that help constrain the model’s predictions. These include well-established principles such as Joule’s law and the conservation of mass, which ensure that the model’s predictions align with physical and scientific laws. Let KKdenote the domain knowledge retrieved from the knowledge base.
@@
To simulate the expert use of historical experiences, AlphaCast constructs a case library by associating each training sample $\langle X_{i},Y_{i}\rangle$ with its optimal candidate model $M_{i}$ selected from the candidate pool $\mathcal{M}$ (detailed in Appendix B.2). To enable efficient retrieval and generation of case-based predictions, we perform K-means clustering (detailed in Appendix A.2) on all $X_{i}$, grouping them into multiple clusters. Each cluster $C_{m}$ contains training samples $X_{i}$ that are most similar to one another. The center $c_{m}$ of each cluster $C_{m}$ is determined by averaging all the samples within the cluster, where the center is the mean of the samples $X_{i}$ in $C_{m}$, calculated as $c_{m}=\frac{1}{|C_{m}|}\sum_{X_{i}\in C_{m}}X_{i}$.
@@ 3.4 The Investigator: Context Preparation
Accurate forecasting requires a well-defined prediction task, a proper understanding of data semantics, and the availability of relevant contextual information. Therefore, AlphaCast constructs a standardized contextual input $I_{\text{in}}$ that integrates time series with textual descriptions. Specifically, $I_{\text{in}}$ comprises the task prompt, the data profile (detailed meanings of the variables), original timestamp information (Wang et al., 2024a; Zeng et al., 2024), and the raw time series $(X_{\text{ex}},X_{\text{en}})$ in original space. This standardized input serves as the foundation for subsequent analysis.
@@ 3.5 The Generator: Reasoning-based Forecasting
Given the consolidated context $I_{\text{con}}$, AlphaCast derives an intermediate prediction $Y_{\text{int}}$ through reasoning-driven generation instead of single-pass regression. The model initializes with the case-based prediction $Y_{\text{case}}$ as a reliable reference. To handle distribution shifts between neighbor $X_{\text{nb}}$ and the current sequence $X_{\text{en}}$, AlphaCast uses the neighbor trajectory $Y_{\text{nb}}$ to adjust for scale and fluctuation differences (Liu et al., 2025c). The framework also incorporates temporal features from the selected feature set $F_{\text{sel}}$, along with domain knowledge $K$ and supplementary context $S$. These elements support more sophisticated adjustments, such as shifting or reshaping, that allow the model to adapt to subtle but important changes in the data, which may be overlooked in simpler models. By incorporating timestamps $T$ for time-aware calibration (e.g., distinguishing weekday variations), the reasoning process is formalized as $Y_{\text{int}}=\mathcal{G}(I_{\text{con}})$, where $\mathcal{G}$ denotes the generator that integrates this context to output the adjusted $Y_{\text{int}}$ for the subsequent stage.
@@ 3.6 The Reflector: Iterative Refinement
In real-world forecasting, predictions are usually reviewed and refined before being finalized. Accordingly, AlphaCast includes a reflection stage to verify the prediction and support iterative refinement. This stage improves the overall reliability and reproducibility of the framework and consists of two parts: (1) prediction evaluation, which assesses whether the prediction satisfies the task requirements and is supported by the available evidence, ensuring that the prediction is reasonable and reliable; and (2) reasoning-trace evaluation, which examines the generated reasoning trace to ensure that each step is valid and consistent, and identifies potential issues such as unsupported claims or logical gaps. Together, these checks help ensure that both the forecast and its underlying reasoning adhere to the task requirements. If issues are found, AlphaCast applies targeted, evidence-supported corrections to $Y_{\text{int}}$ to obtain a refined prediction. The refined result is then re-evaluated with the same criteria. If acceptable, AlphaCast outputs the forecast; otherwise, it returns feedback to the previous stage for another update, repeating until an acceptable result is obtained.
@@ 4 Experiments
In this section, we conduct experiments on various forecasting benchmarks comparing AlphaCast with mainstream approaches to demonstrate its effectiveness.
@@ 4.2 Main Results
Table 1 compares AlphaCast against twelve baselines across short-term and long-term benchmarks. Overall, AlphaCast achieves state-of-the-art performance with superior accuracy across most metrics. In short-term scenarios, AlphaCast shows outstanding stability in the face of high volatility, effectively managing unpredictable fluctuations. Unlike statistical methods such as ARIMA, which struggle with rapid non-linear fluctuations and inherent stochastic volatility in price dynamics, or deep learning models such as PatchTST, which may over-smooth dynamics, AlphaCast effectively captures sharp peaks by leveraging historical references to avoid phase shifts. This gives it a significant advantage on datasets like BE and FR. In long-term settings, AlphaCast exhibits superior forecasting capability, especially on datasets like Windy Power and Sunny Power, which exhibit complex environmental dependencies. While foundation models such as Chronos perform competitively due to their zero-shot generalization, AlphaCast maintains a consistent lead through its reasoning and refinement capabilities that enforce causal constraints derived from diverse contexts. For instance, on the Windy Power dataset, AlphaCast substantially outperforms baselines like DLinear. This superiority across different scales validates the effectiveness of the interactive, multi-turn reasoning process in handling diverse temporal patterns and ensuring forecasting stability.
@@ 4.3 Ablation studies
This section validates the AlphaCast framework through a multi-perspective ablation study, covering both architectural stages, contextual input, and auxiliary toolkit modules.
@@
To validate the workflow, we conduct an ablation study by removing the toolkit, reasoning, and reflexion. As shown in Table 2, the full model consistently achieves best performance, validating component collaboration. Specifically, removing the toolkit yields highest errors in scenarios like BE and ETTm, confirming external context is foundational. Furthermore, comparing the baseline lacking reasoning, which serves as the state-of-the-art candidate model backbone, with the full model reveals that reasoning-based forecasting substantially improves accuracy, evidenced by notable error reduction on the FR dataset. Most importantly, results on the Sunny Power dataset highlight the indispensability of the reflexion module, where the variant excluding self-correction performs worse than even the non-reasoning baseline. This anomaly suggests that without self-correction, generative reasoning may be prone to hallucinations or overfitting, an issue rectified by the full model acting as a safeguard for reliability and physical consistency.
@@
To verify the necessity of constructing a comprehensive context $I_{\text{con}}$, we conduct a component-wise ablation study. As shown in Figure 3, the complete AlphaCast framework consistently yields the lowest MAE, confirming the synergistic contribution of all elements. Specifically, removing timestamps causes significant degradation on periodic datasets such as PJM and Sunny Power, highlighting the critical role of explicit temporal identifiers in capturing seasonality that numerical sequences alone cannot convey. Similarly, the exclusion of exogenous variables and static data profile leads to substantial error increases on tasks like ETTm and NP, demonstrating that external drivers and domain-specific inductive biases are essential for reasoning about complex system behaviors beyond historical autoregressive patterns.